1) A daring 51 kg swimmer dives off a cliff with a running horizontal leap.

A)What must her minimum speed be just as she leaves the top of the cliff so that she will miss the ledge at the bottom, which is L = 2.40 m wide and h = 9.40 m below the top of the cliff?

2)A physics book slides off a horizontal tabletop with a speed of 3.40 m/s. It strikes the floor in 0.400 s. Ignore air resistance.

A)Find the height of the tabletop above the floor. Express your answer in meters.

B) Find the horizontal distance from the edge of the table to the point where the book strikes the floor. Express your answer in meters.

C) Find the horizontal component of the book's velocity, just before the book reaches the floor. Express your answer in meters per second.

D) Find the vertical component of the book's velocity just before the book reaches the floor. Express your answer in meters per second.

E) Find the magnitude of book's velocity, just before the book reaches the floor. Express your answer in meters per second.

F) Find the direction of book's velocity, just before the book reaches the floor. Express your answer in degrees.

3)A wall clock has a second hand 23.0 cm long. What is the radial acceleration of the tip of this hand? Express your answer in centimeters per second squared.
Answer #1

1)

Using 2nd equation of motion,

S = ut + 0.5 at^2

9.4 = 0 + 0.5 x 9.8 x t^2

Time, t = 1.39 sec

Distance, d = ut

Velocity, v = 2.40/1.39 = 1.73 m/s

Comment in case any doubt please rate my answer...

Please post other questions separately...